Give an example of an series Σan such that Σan is not convergent but the sequence (an) converges to 0


1
Expert's answer
2021-05-07T11:52:28-0400

Solution:

Assume "a_n=\dfrac{\ln (n)}{n}"

"\lim _{n \rightarrow \infty} \frac{\ln n}{n}"

"\\=\lim _{n \rightarrow \infty} \frac{1 / n}{1}" [Using L' Hopital rule]

"=0"

Next, "\Sigma a_n= \sum _{n=1}^{\infty \:}\frac{\ln \left(n\right)}{n}"

Consider "\int_{1}^{\infty} a_{n} d n=\int_{1}^{\infty} \frac{\ln n}{n} d n=\left[\frac{(\ln n)^{2}}{2}\right]_{1}^{\infty}= diverges"

So, by integral series test, "\int_{1}^{\infty} a_{n} d n" is divergent, then so is "\sum_{n=1}^{\infty} a_{n}"

Thus, we have "\sum_{n=1}^{\infty} \frac{\ln n}{n}" is divergent.